# point-to-point OTG Online Trajectory Generation

### v1

@@ IV-C Optimization
Inverse kinematics optimization aiming for global optimality is significant computationally expensive. Recent study [6] reports an average computation time of 77.6 seconds to optimize a system with 10 DoF using the similar objective functions in IV-B. The optimizing problems of a dual-arm system involves commonly 12-14 DoF, resulting in an expanded optimization space. Due to the curse of dimensionality, this increase significantly elevates computational expense, rendering such methods impractical.
@@
The other non-convex optimization algorithms such as [28], trade global optimality for potentially faster computation. To improve the chances of finding a better solution, a parallel optimization approach is often employed [29]. This method uses multiple initial guesses in parallel, and the best result is selected as the optimal solution. Since the mapping relationship between Cartesian coordinates and the joints of a redundant manipulator is not unique, the number and distribution of initial guesses can significantly influence the final optimization result.
The entire objective function is
$f(\mathbf{q})=w_{p}\cdot\chi_{p}(\mathbf{q})+w_{o}\cdot\chi_{o}(\mathbf{q})+w_{m}t\cdot\chi_{\text{mt}}(\mathbf{q},\mathbf{q}_{0})+w_{b}\cdot\chi_{b}(\mathbf{q})$ (9)
@@
where $w_{i}$ represent the weight for the corresponding objective terms, and $\chi_{b}(\mathbf{q})$ is the joint limit cost proposed in [29]. Initially, Halton sampling is employed to generate samples from the configuration space that are free of collisions, serving as initial guesses for the optimization process. The step directions for the optimization are then computed based on the gradient of the defined objective function.
Once the step direction is determined, a line search is conducted. This involves scaling the step direction with a discrete set of magnitudes [29]. The optimal magnitude from this set is selected using the Armijo and Wolfe conditions, ensuring that the step size is sufficient for descent and satisfies the curvature condition, which is crucial for avoiding overshooting and ensuring stable convergence in IK problems.
@@
Typically, the IK problem defines the optimal solution as the one that minimizes the pose error. Since our method also consider the motion time, the metric for selecting the optimal solution can be defined in several ways: it can prioritize the pose error, the motion time, or a combination of both. This flexibility allows the optimization to balance accuracy in achieving the desired pose with the efficiency of motion, depending on the specific requirements of the task.
IV-D Execution Time Approximator
Using surrogate loss such as distance between joint configurations as optimization objectives assumes that resultant point-to-point motion between the start and target configuration is collision-free and executable. For a dual-arm setting, this assumption is usually not legitimate and therefore a post-processing step such as path planning is needed. This additional step introduce an gab between the surrogate loss and the actual execution time. Directly coping execution time in the optimization does not has such problem.
For dual arms mounted on a fixed base, we can deploy offline optimal path planning or trajectory optimization algorithms to compute the collision-free motion between two arbitrary joint configuration and at the end acquire the actual execution time. However, this paradigm can not be included into the multi-objective optimization due to its computation complexity. Instead, we collect data from these offline methods and use a multilayer perceptron (MLP) $f_{\theta}$ to approximate the execution time during the optimization. Previous works [30] show that MLPs are capable to capture the latent relation between joint configurations and time. Furthermore, MLPs can not only achieve fast inference time during optimization, but also support batch operations and provide gradients for the optimization. In practice, we use positional encoding $\mathbf{q}_{PE}$ = [$\mathbf{q}$, $\cos\mathbf{q}$, $\sin\mathbf{q}$] as input for the approximator. Given the start and target configuration $\mathbf{q}_{0}$, $\mathbf{q}_{t}$, the approximator outputs
$\hat{t}=f_{\theta}(\mathbf{q}_{PE,0},\mathbf{q}_{PE,t}).$ (10)
Literature [31][32] suggests that this can improve the expressiveness of the MLP.
@@ V Experimental Evaluations
The experimental setup for the dual-arm system, shown in 3, consists of a UR5 and a KUKA LBR iiwa 14 R820 robot, with a combined total of 13 DoF. The KUKA iiwa is designated for object grasping due to its higher load capacity, while a high-resolution 3D lidar Robin W from Seyond is mounted on the UR5’s end-effector for scanning purposes.
V-A Training Dataset
Given a start and target configuration, the time approximator provides an estimated execution time with implicit consideration of potential collision risk. To train the time approximator, we generate 250,000 collision-free start and target pairs. Then, we feed these pairs to the TOPPRA [18] trajectory generator and Curobo TrajOpt [29] to collect respective execution time. The difference between these two trajectory generators is that, the one from Curobo takes collision into account and always returns collision-free trajectories while the TG methods can return trajectories that are not executable. In our generated dataset, approximately 50% of the trajectories generated by TOPPRA are not collision-free and, therefore, not executable in practice.
V-B Simulation
V-B1 Evaluation Simulation Setup
In the simulation, a general evaluation of the proposed approach is conducted. This involves generating a random collision-free initial joint configuration $\mathbf{q}_{0}$ and a random relative Cartesian pose $\mathbf{p}_{R,t}$. A random reference target joint configuration $\mathbf{q}_{t}$ is sampled, and the absolute TCP poses of the two robots are then calculated individually.
For comparative analysis, three baselines are provided. These baselines utilize the random reference $\mathbf{q}_{t}$ directly as the endpoint, perform optimization on Eq. (8) based on absolute and relative Cartesian pose as motion time term in the objective function. The resulting motion time based on the IK solutions is subsequently calculated using the TOPPRA algorithm.
@@ V-B1 Evaluation Simulation Setup
The robot joint limits used when generating the dataset for the network align with those in TOPPRA, including velocity $\left\{3.15,3.15,3.15,3.2,3.2,3.2\right\}$ and $\left\{10.0,10.0,10.0,10.0,10.0,10.0,10.0\right\}$ $\frac{rad}{s}$, acceleration $\left\{5.0,5.0,3.0,2.0,2.0,2.0\right\},\left\{5.0,5.0,3.0,2.0,2.0,2.0,2.0\right\}$ $\frac{rad}{s^{2}}$, and jerk constraints for UR5 and IIWA respectively $\left\{500.0,500.0,500.0,500.0,500.0,500.0\right\}$, $\left\{500.0,500.0,500.0,500.0,500.0,500.0,500.0\right\}$ $\frac{rad}{s^{3}}$. All optimization parameters are summarized in Tab. I. To explore the optimal solution, the number of initial guesses and optimization iterations are set to a high value, namely $4096$ initial guesses and $500$ iterations. This ensures a comprehensive evaluation of the solution space. The weights for position and orientation error are $2000,2500$ and the execution time weights are set as $250,500$ for joint distance and execution time approximator, respectively.
Metrics Ref. $\mathbf{q}_{t}$ (A) (B) (C) (D) (E) (F)
execution time by TOPPRA 2.71 2.618 2.363 2.477 2.053 2.450 2.807
Position error 0 0.0004 0.0002 0.0022 0.0016 0.0003 0.0003
IK success rate - 100 100 98 100 100 100
TABLE I: Resulting execution time with random sampling from 100 IK instances (A) Given TCP-pose with distance loss (B) Given TCP-pose with trained model and best time selection (C) Relative pose with distance loss (D) relative pose with trained model and best time selection (E) relative pose with trained model and best cost selection (F) relative pose with trained model and best pose error
@@
Figure 4: Execution time comparison between reference, method (C) and (G) for collision-free trajectory generation: The starting pose and the reference target shown are randomly generated. The starting joint configuration, reference target configuration, and the IK solution generated by method (C) and (G) in Tab. II are illustrated with the corresponding position and velocity profile over time generated by TrajOpt.
Metrics Ref. $\mathbf{q}_{t}$ (C) (G)
execution time by TOPPRA 2.483 2.4529 2.394
execution time by TrajOpt 3.522 3.513 3.40
position error 0 0.0016 0.002
IK success rate - 100 100
@@
TABLE II: Resulting execution time with task related relative pose from 100 IK instances (C) Relative pose with distance loss (G) relative pose with trained collision-free model
V-B2 Execution Time Evaluation
@@
The execution time results indicate that using absolute Cartesian coordinates as position targets is less effective than relative positions for execution time optimization. The distance loss (C in the Tab. I) and the proposed execution time approximator with cost based IK solution selection (E) demonstrate comparable performance in optimizing the execution time. It is worth noting that due to the greater optimization space, the positional error of (E) is significantly smaller. If the solution to success is sorted only by execution time cost, the solution obtained by proposed ETA-IK is almost $25\%$ shorter than the execution time of the randomly generated ref configuration, performs best among all comparison methods.
V-B3 Implicit Collision Consideration
@@
For the approximator trained by the collision-free dataset, the NBV candidate poses are utilized as target relative poses. Since these scanned positions are object-centered, they are more prone to collisions, thereby highlighting the influence of the collision-aware approximator on the IK solution. As shown in the Tab. II, the trained approximator demonstrates faster performance in both TG methods, with and without collision checking. Method (C), which employs relative position and distance loss as objectives, results in minimal joint changes (A typical example with random sampled starting and target is shown in Fig. 4). However, due to the need for collision avoidance, the final collision-free trajectory execution time is longer than method (G). The symmetry of the velocity profile indicates the presence of obstacle avoidance behavior. While the reference trajectory is planned without encountering obstacles, resulting in a completely symmetrical velocity profile. The longer joint movement distances increased the execution time. In contrast, the IK solution obtained through the proposed approximator optimization exhibits relatively less obstacle avoidance behavior and achieves optimal execution time among the three methods considered.
@@ V-C Real Robot Experiments
The real robot experiments validated the motion generation pipeline, demonstrating that the generated trajectories perform effectively both in the simulation environment and on real robots. A video of the experiment is attached to provide a visual demonstration of the system’s performance.
@@ VI Discussion and Conclusion
In this work, we presented ETA-IK, an Execution-Time-Aware Inverse Kinematics method for dual-arm systems, with a specific application in scanning unknown objects during the dismantling of a nuclear facility. Our primary focus was on optimizing motion execution time by leveraging the redundancy of dual-arm systems. The results demonstrated that integrating motion execution time directly into the IK optimization framework yields significant improvements in efficiency compared to traditional methods that rely on surrogate metrics, such as joint distance.
The experiments showed that our proposed execution time approximator effectively improves the overall system performance. By incorporating collision avoidance into the optimization process, our approach could account for self-collision without introducing substantial computational overhead. This was particularly evident in both simulated and real-world scenarios, where the system was able to achieve more efficient execution while maintaining the relative pose between the two robots.
In addition, we have identified several limitations in this work. One is the additional inference time required for the neural network-based approximator compared to simple distance-based metrics. However, this drawback is mitigated by the overall improvement in the motion execution time. Future work could focus on reducing the computational overhead of the approximator and exploring the application of this approach to multi-robot systems with higher degrees of freedom.

### v2

@@ IV-D Optimization
Inverse kinematics optimization aimed at achieving global optimality is computationally expensive. A recent study [13] reports an average computation time of 77.6 seconds to optimize a system with 10 DoF using the similar objective functions in IV-B. Due to the curse of dimensionality, applying this method for online IK is impractical.
@@
The other non-convex optimization algorithms such as [14], trade global optimality for potentially faster computation. To improve the chances of finding a better solution, a parallel optimization approach is often employed [28]. This method uses multiple initial guesses in parallel, and the best result is selected as the optimal solution. Since the mapping relationship between Cartesian coordinates and the joints of a redundant manipulator is not unique, the number and distribution of initial guesses can significantly influence the final optimization result. The entire objective function is
$f(\bm{q})=w_{p}\cdot\chi_{p}(\bm{q})+w_{o}\cdot\chi_{o}(\bm{q})+w_{mt}\cdot\chi_{\text{mt}}(\bm{q},\bm{q}_{0})+w_{b}\cdot\chi_{b}(\bm{q})$ (10)
@@
where $w_{i}$ represent the weight for the corresponding objective terms, and $\chi_{b}(\bm{q})$ is the joint limit cost proposed in [28]. Initially, Halton sampling is employed to generate samples from the configuration space that are free of collisions, serving as initial guesses for the optimization process. The step directions for the optimization are then computed based on the gradient of the defined objective function.
@@
Typically, the IK problem defines the optimal solution as the one that minimizes the pose error. Since the proposed method also consider the motion time, the metric for selecting the optimal solution can be defined in several ways: it can prioritize the pose error, the motion time, or a combination of both. This flexibility allows the optimization to balance accuracy in achieving the desired pose with the efficiency of motion, depending on the specific requirements of the task.
@@ V Experimental Evaluations
Pick-IK-1 Pick-IK-256 ETA-IK-20 ETA-IK-2k ETA-IK-200k
Surrogate Dist. Est. Time Surrogate Dist. Est. Time Est. Time
Execution Time [s] 2.347 - 2.073 2.055 1.963 1.926 1.892
Joint Configuration Dist. $rad^{2}$] 5.50 - 4.71 4.73 4.33 4.33 4.28
Computation Time [ms] 15.78 833.46 59.40 95.77 62.48 99.70 980.23
IK Success Rate [%] 99.5 0 100 100 100 100 100
TABLE I: Evaluation results based on random sampling from 10,000 IK instances. Two variants of optimization objectives for ETA-IK are evaluated. While both of the variants use relative pose, Surrogate Dist. refers to the variant using joint configuration distance loss in the objective function and Est. Time indicates the variant directly coping with the trained time approximator. Bold values highlight that, even with similar joint distances, the execution time approximator identifies more optimal solutions. Italic values show that increasing the number of optimization seeds further improves execution time. Underlined values indicate that multi-threaded GPU inference ensures scalability without increasing computation time.
V-A General Evaluation in Simulation
Initially, we configured a simple simulation environment featuring a dual Franka setup without environmental collisions to evaluate the general performance of the proposed method. This setup involves generating a random, self-collision-free initial joint configuration $\bm{q}_{0}$ and a random relative Cartesian pose $\bm{x}_{R,target}$.
For comparative analysis, we use Pick-IK (Reimplementation of Bio-IK [17] by MoveIt2) as the baseline. The optimization includes joint configuration distance as an additional metric. We also evaluate multi-threaded computation on the CPU, which is provided as a built-in functionality of Pick-IK. The evaluation criteria include:
  • The trajectory execution time after post-processing with TOPPRA
  • The difference between the initial joint configuration and the computed solution
  • The computation time required to solve the IK problems
  • The success rate of solving the IK problems
For Pick-IK, the default setup with a single seed and CPU multi-threading with 256 seeds is selected as the baseline. The global mode is used for the PTP IK problem, and the IK solver timeout is set to $0.5$ seconds. The weight for the joint displacement is set to $0.002$, as higher values result in an extremely high failure rate.
For ETA-IK, the default parameter values of Curobo [28] are used. The joint configuration distance weights are set to $5$ and $20$ for the MPPI and Newton methods. The learned approximator applies displacement weights of $5$ and $50$, respectively.
The experiments were conducted on a desktop PC with an Intel i9 processor and an NVIDIA RTX 4090 GPU. The execution time approximator was also trained locally on this machine, requiring approximately 1 hour. Statistical results were evaluated over 10,000 independent inverse kinematics queries. Depending on the trajectory generation method used, the dataset generation time for 250,000 trajectories ranged from several hours to approximately one day.
V-A1 Execution Time Evaluation
In general, IK solutions obtained from the parallel multi-objective optimization framework result in shorter execution times. Even when the joint differences are identical, the execution time approximator can identify a more efficient solution for execution (see bold results in Table I). Additionally, increasing the number of optimization seeds further improves the solution quality (see italicized results in Table I).
V-A2 Computation Time Evaluation
Although the proposed method generally exhibits longer computation times compared to Pick-IK, and the execution time approximator incurs additional inference overhead, it is important to note that the computation time remains unchanged when using 20 or 2000 seeds due to the multi-threading capabilities of the GPU (see underlined results in Table I). In contrast, Pick-IK with 256 threads experiences a significant computational overhead compared to its single-threaded counterpart. Moreover, the success rate of Pick-IK with 256 threads is zero, as all computations exceed the predefined time limit.
While the initial evaluation benchmarks the proposed method against baseline methods and standard metrics, the subsequent ablation study investigates how different design choices within the proposed framework affect performance, particularly in terms of shorter execution times and implicit collision handling.
V-B Ablation Study
@@
The real experimental setup for the dual-arm system, shown in Fig. 2, consists of a UR5 and a KUKA LBR iiwa 14 R820 robot, with a combined total of 13 DoF. The KUKA iiwa is designated for object grasping due to its higher load capacity, while a high-resolution 3D LiDAR Robin W from Seyond is mounted on the UR5’s EE for scanning purposes.
In the simulation, additional static obstacles, such as the table, LiDAR, and viewpoint sphere, are present compared to the environment described in Section V-A (as shown in Fig. 5). To implement absolute TCP pose IK, we sample a random, collision-free initial joint configuration $\bm{q}_{0}$ and a random reference target joint configuration $\bm{q}_{target}$. The relative Cartesian pose $\bm{x}_{R,target}$ and the absolute TCP poses $\bm{x}_{A,target},\bm{x}_{B,target}$ of the two robots are then computed individually.
@@
The robot joint limits used when generating the dataset for the network align with those in TOPPRA, including velocity $\left\{3.15,3.15,3.15,3.2,3.2,3.2\right\}$ and $\left\{10.0,10.0,10.0,10.0,10.0,10.0,10.0\right\}$ $\frac{rad}{s}$, acceleration $\left\{5.0,5.0,3.0,2.0,2.0,2.0\right\},\left\{5.0,5.0,3.0,2.0,2.0,2.0,2.0\right\}$ $\frac{rad}{s^{2}}$ constraints for UR5 and IIWA respectively. The maximum jerk for all the joints is $500\frac{rad}{s^{3}}$.
For the ablation study, different design choices are varied for evaluation. Methods (A) and (C) optimize Eq. (8), using absolute and relative Cartesian poses as the surrogate execution time term in the objective function. Method (B) employs the trained execution time approximator but considers absolute Cartesian poses. Methods (D), (E), and (F) use the trained execution time approximator with different solution selection criteria, namely best time, best cost, and best pose error. All optimization parameters are summarized in Tab. II. To explore the optimal solution, the number of initial guesses and optimization iterations are set at a relatively high value, namely $4096$ initial guesses and $120$ iterations. This ensures a comprehensive evaluation of the solution space while maintaining computation time within a practically acceptable range. The weights for position and orientation error are $2000,2500$ and the execution time weights are set to be $250$ for joint distance and $500$ for execution time approximator.
Exec. T. Error SR
[s] [$m$] [%]
Reference $\bm{q}_{t}$ 2.71 0 -
(A) Abs. TCP + Surrogate Dist. 2.618 0.4 100
(B) Abs. TCP + Est. Time + Best Time 2.363 0.2 100
(C) Relative + Surrogate Dist. 2.477 2.2 98
(D) Relative + Est. Time + Best Time 2.053 1.6 100
(E) Relative + Est. Time + Best Cost 2.450 0.3 100
(F) Relative + Est. Time + Best Pose 2.807 0.3 100
TABLE II: Ablation study: execution time with random sampling from 100 IK instances using different input, objective functions and selection criteria. Execution time, position error and success rate (SR) are used as metrics. Specifically, absolute TCP and relative pose are two options for input desired pose. The objective options are marked in blue and the selection criteria are marked in purple.
@@
Figure 5: Execution time comparison between reference, method (C) and (G) for collision-free trajectory generation: The starting pose and the reference target shown are randomly generated. The starting joint configuration, reference target configuration, and the IK solution generated by method (C) and (G) in Tab. III are illustrated with the corresponding position and velocity profile over time generated by TrajOpt.
Exec. T. Exec. T. Error SR
(TOPPRA) (TrajOpt)
[s] [s] [$m$] [%]
Reference $\bm{q}_{t}$ 2.483 3.522 0 -
(C) Rel. + Sur. Dist. 2.4529 3.513 1.6 100
(G) Rel. + Est. T. (Col. free) 2.394 3.40 2 100
@@
TABLE III: Resulting execution time with task-related relative pose from 100 IK instances. (C) Relative pose with distance loss; (G) relative pose with trained collision-free model.
V-B1 Execution Time Evaluation
@@
The execution time results indicate that using absolute Cartesian coordinates as position targets is less effective than relative positions for execution time optimization. The distance loss (C in the Tab. II) and the proposed execution time approximator with cost based IK solution selection (E) demonstrate comparable performance in optimizing the execution time. It is worth noting that due to the greater optimization space, the positional error of (E) is significantly smaller. If the solution to success is sorted only by execution time cost, the solution obtained by proposed ETA-IK is almost $25\%$ shorter than the execution time of the randomly generated ref configuration, performs best among all comparison methods.
V-B2 Implicit Collision Consideration
@@
For the approximator trained by the collision-free dataset, the NBV candidate poses are utilized as target relative poses. Since these scanned positions are object-centered, they are more prone to collisions, thereby highlighting the influence of the collision-aware approximator on the IK solution. As shown in the Tab. III, the approximator trained with the collision-free dataset (G) demonstrates faster performance in both TG methods, with and without collision checking. Method (C), which employs relative position and distance loss as objectives, results in minimal joint changes (A typical example with random sampled starting and target is shown in Fig. 5). However, due to the need for collision avoidance, the final collision-free trajectory execution time is longer than method (G). The symmetry of the velocity profile indicates the presence of obstacle avoidance behavior. While the reference trajectory is planned without encountering obstacles, resulting in a completely symmetrical velocity profile. The longer joint movement distances increased the execution time. In contrast, the IK solution obtained through the proposed approximator optimization exhibits relatively less obstacle avoidance behavior and achieves optimal execution time among the three methods considered.
@@ V-C Real Robot Experiments
The real robot experiments validated the motion generation pipeline, demonstrating that the generated trajectories perform effectively both in the simulation environment and on real robots. Videos demonstrating the system’s performance are available on the https://yucheng-tang.github.io/eta-ik-website/.
@@ VI Discussion and Conclusion
In this work, we presented ETA-IK, an Execution-Time-Aware Inverse Kinematics method for dual-arm systems, with a specific application in scanning unknown objects during the dismantling of a nuclear facility. Our primary focus was on optimizing motion execution time by leveraging the redundancy of dual-arm systems. Unlike traditional trajectory generation techniques or physics-based simulators, ETA-IK leverages a trained execution time approximator that can be efficiently integrated into the optimization framework. This is enabled by its ability to support parallel inference and to provide differentiable outputs.
Experimental results demonstrated that incorporating motion execution time directly into the IK optimization framework yields significant improvements in efficiency compared to traditional methods that rely on surrogate metrics, such as joint configuration distance. By incorporating collision avoidance into the optimization process, our approach could account for self- and static collision without introducing substantial computational overhead. This was particularly evident in both simulated and real-world scenarios, where the system was able to achieve more efficient execution while maintaining the relative pose between the two robots.
In addition, we identified a limitation of the proposed method: the additional inference time required by the neural network-based execution time approximator compared to simple distance-based surrogate metrics. However, this computational overhead remains nearly constant as long as the number of parallel threads does not exceed the capacity of the GPU. Nevertheless, it should be acknowledged that the proposed approach relies on access to a sufficiently powerful GPU to support efficient parallel inference and optimization within practical time constraints. Future research may explore reducing the overhead of the approximator further and learning more effective initial pose distributions—potentially replacing Halton sampling—with methods such as the one proposed in [31].